• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Holmstrom

    Log entry time 00:01:49 on July29,2003

    Entry number 107795

    keyword=Swing Shift Summary

    16:00 Begining shift, the previous shift finsihed a harp scan, but the OTR looks strange, MCC has been called. We are waiting for the OTR to give a better result, and then we will run spot++ and start kinimatic 5.14.

    16:20 Olivier talked to MCC and estimate for beam is 30 minutes.

    16:22 J.P. would like one more Harp Scan, to improve from 80x136.

    16:50 Beam returns. I have asked for them to widen the beam in x, and then we will do another Harp scan. They are going to try to test the fast feedback, which will take about 5 minutes.

    17:10 Beam is through the Compton, and they are trying adjust the beam width in x.

    17:12 MCC called and they will start the 5 minute fast feedback study now.

    17:34 Harp Scan 2267 - #5 159x112 #6 82x158
    17:40 Harp Scan 2268 - #5 166x103 #6 88x147
    17:51 Harp Scan 2269 - #5 170x107 #6 98x141

    17:53 Compton studies underway.

    18:05 Asked for raster on.
    18:07 We will do a compton measurement and than a moller after.
    18:15 We are trying to fix the raster size.

    18:24 To get 3.5x3.5 raster we request 15x4 from MCC.

    18:25 Requset the beam masked, to move target to empty reference cell position.

    18:30 Began Kinimatic 5.14

    Kin. 5.14
    Right arm momentum = 1.482, angle=12.5 deg.
    Left arm momentum = 3.17, angle=18 deg.
    Septum Current = 137 A

    Raster on
    Beam Current = ?.0 amp

    Right arm run
    Time Run Num. Target PS1 PS2 beam lamba/2 target lamba/2 Events
    18.28-18:36 2956 Reference Empty 1 1 IN OUT 1.02M
    18.45-18:55 2958 Reference N2 10 9 IN OUT 1.85M
    22.08-22:22 2961 3He(0 deg) 1 1 IN OUT 3.2M
    22.24-22:38 2962 3He(0 deg) 1 1 IN OUT 3M
    22.39-22:44 2963 3He(0 deg) 1 1 IN OUT 1.1M
    22.53-22:07 2964 3He(0 deg) 1 1 OUT OUT 3M
    23.08-23:21 2965 3He(0 deg) 1 1 OUT OUT 3M
    23.22-23:29 2966 3He(0 deg) 1 1 OUT OUT 1.3M
    23.39-23:?? 2967 3He(270 deg) 2 2 OUT OUT 3M
    23.39-23:?? 2968 3He(270 deg) 2 2 OUT OUT ?M


    18:56 Request beam off and target mask from MCC, so we can move to empty target position for a Moller measurement.

    21:40 Moller finished, waiting. Bleedthrough during Moller of 2.0%.
    21:45 We checked spot++ and we think that the x has narrowed. We ask for 17x4 in the rastor.
    21:51 We checked spot++ and we get 3.5x3.5 with a rastor request of 17x4.

    21:52 Request MCC remove beam and maks the target, so that we can move to 3He position for NMR and analysis. (Note: 3He and Pickup coil position are close enough that Vince ok'd using the 3He position.)

    22:00 NMR result P=39.28%.

    22:26 Changed the stop point of the right HRS DAQ to 3M from 3.2M.

    22:45 Request MCC take out the beam 1/2 wave plate. David went to talk to Hall B.

    22:55 We had a period of about 40 minutes where Hallog was not updating. We have also been getting error messages at start of runs see hallog entry:

    23:03 BTA software is not working properly with 1.5uA. We are getting a lot of BNA, which is actuall ABU.

    23:30 We did not correct for the cuts on data, so we will correct for that with the kin. 5.14 270 deg. data.

    23:31 Olivier would like beam current above 2uA, so we will ask for more beam and prescale the next set of data.

    23:35 We have rotated the target 270 deg, and we have asked for 3uA, and we will prescale the triggers by 2.

    00:00 Shift over we are almost finished with the 5.14 kinimatic.